OpenStudy (aryana_maria2323):

Please help. Will FAN AND MEDAL! A firework is launched at the rate of 10 feet per second from a point on the ground 50 feet from an observer. To 2 decimal places in radians per second, find the rate of change of the angle of elevation when the firework is 40 feet above the ground.

OpenStudy (aryana_maria2323):

can you please help @Photon336

OpenStudy (aryana_maria2323):

Where did the graph go that you just put up?

OpenStudy (aryana_maria2323):

Okay I don't understand?

OpenStudy (photon336):

I'm guessing we could find a way to express theta the angle of elevation as a function of time. but we know that from our physics equations that the change in the vertical distance is equal to \[V_{o}Sin(\theta)t+\frac{ 1 }{ 2 }at^{2} = {y-y_{0}}\] while the change in horizontal distance is equal to this. \[V_{o}Cos(\theta)t = (x-x_0)\] to find theta we know that if we know what the y distance traveled over x right and we take the inverse tangent we can find the angle theta. \[\theta = tangent^{-1}\frac{ y }{ x }\] \[\theta = tangent^{-1}(\frac{ y-y_{0}}{ x-x_{0} })\] \[\theta = tangent^{-1}\frac{ V_{o}Sin(\theta)+\frac{ 1 }{ 2 }at^{2} }{ V_{o}*Cos(\theta)t }\] \[\frac{ d~\theta }{ dt} = \frac{ d~\theta }{ dt } (tangent^{-1}\frac{ \cancel\V_{o}Sin(\theta)+\frac{ 1 }{ 2 }at^{2} }{ \cancel\V_{o}*Cos(\theta)t })\] omg this is really nasty \[\frac{ d~\theta }{ dt} = \frac{ d~\theta }{ dt } (tangent^{-1}\frac{ Sin(\theta)+\frac{ 1 }{ 2 }at^{2} }{ Cos(\theta)t })\] @Kainui please help lol i'm lost.

OpenStudy (photon336):

@ganeshie8

ganeshie8 (ganeshie8):

I think we should assume that the fire cracker is launched vertically up

ganeshie8 (ganeshie8):

But then the firework won't go that high. Something seems to be wrong with the question here...

OpenStudy (kainui):

I think this is not a well thought out calculus problem where gravity doesn't exist